OpenStudy (anonymous):

8.37 x 10^8/2.7 x 10^3 Hey I need some help understanding the steps to solving this. I'm so lost.

OpenStudy (gibbs):

1 Simplify 103 to 1000 8.37x×1082.7x×1000 2 Use Quotient Rule: xaxb=xa−b 3100x1+1×108 3 Simplify 1+1 to 2 3100x2×108 4 Regroup terms 3100×108x2 is the answer

OpenStudy (gibbs):

do you get it?

OpenStudy (anonymous):

Um not exactly :(

OpenStudy (anonymous):

@gibbs can you explain it better?

OpenStudy (gibbs):

Simplify 103 to 1000

OpenStudy (gibbs):

Simplify 103 to 1000 witch you will get 8.37x×1082.7x×1000

OpenStudy (gibbs):

then the rest should be self explained if you need help anymore im not your guy srry

OpenStudy (anonymous):

I'm dividing one side by the other though.

OpenStudy (gibbs):

Simplify 10^3 to 1000 thats what wasnt right hang on

OpenStudy (gibbs):

Simplify 10^3 to 1000 8.37x×10^8 2.7x *1000 2 Use Quotient Rule: xaxb=xa−b 3100x^1^+^1×10^8 3 Simplify 1+1 to 2 3100x^2×10^8 4 Regroup terms 3100×10^8x^2

OpenStudy (gibbs):

ok now does that make sense?

OpenStudy (anonymous):

I just figured it out. So the way my textbook is saying take 8.37/2.7 which is 3.1. Then 10^8/10^3, it becomes 10^5 because you subtract the exponents. so it's 3.1 x 10^5 which equals 310,000.